OpenStudy (anonymous):

what is 5/18 + (-5/6)? A. -5/9 B. -1/12 C. 3/8 D. 1/8

TheSmartOne (thesmartone):

multiply 5/6 by 3 on both the numerator and denominator :)

OpenStudy (anonymous):

15/ 18 @TheSmartOne

TheSmartOne (thesmartone):

ok now we have: 5/18 - 15/18 what is 5 -15 =?

OpenStudy (anonymous):

5 - 15 = -10 15 - 5 = 10

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

5/18 - 15/18 = (5-15)/18 = -10/18

TheSmartOne (thesmartone):

divide both the numerator and denominator by 2

OpenStudy (anonymous):

-5/9

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

coorect, good job :)